OpenStudy (freckles):

\[\text{ Let } \\ u=1+\frac{x^3}{3!}+\frac{x^6}{6!}+\frac{x^9}{9!}+\cdots \\ v=x+\frac{x^4}{4!}+\frac{x^7}{7!}+\frac{x^{10}}{10!}+\cdots \\ w=\frac{x^2}{2!}+\frac{x^5}{5!}+\frac{x^8}{8!}+\cdots \\ \text{ Show that } u^3+v^3+w^3-3uvw=1.\] \[

OpenStudy (kainui):

First idea: \[\Large w'' = v' = u \\ \Large u+v+w = e^x\]

OpenStudy (kainui):

I guess also u' = w and others but yeah, I don't know if that's useful to us or not.

OpenStudy (perl):

you can factor the left side of that equation

OpenStudy (kainui):

good idea, let's take (u+v+w)^3 and see what terms we can collect and throw around, maybe that helps too

OpenStudy (perl):

(w+u+v)*(u^2-u*v-u*w+v^2-v*w+w^2)= 1

OpenStudy (freckles):

\[e^x(u^2+w^2+v^2-uv-uw-vw)\]

OpenStudy (perl):

if you can show the latter term is equal to e^-x , you are done

OpenStudy (perl):

but this could be a blind alley :)

OpenStudy (freckles):

\[(u+v+w)^3 -3u^2v-3u^2w-3uv^2-6uwv-3uw^2-3v^2w-3vw^2\] just seeing what the other way looks like I think I got rid of all the trash there

OpenStudy (perl):

right that would be a better approach

OpenStudy (freckles):

i have no clue what the better approach is :p

OpenStudy (freckles):

maybe we can factor your other factor there @perl

OpenStudy (freckles):

i don't have you mad fancy factoring skills though

OpenStudy (freckles):

how did you factor that one thingy or how did you see that

OpenStudy (perl):

I used maple , a computer algebra system

OpenStudy (freckles):

ah i was trying to get wolfram to factor it for me but it told me to get lost

OpenStudy (perl):

:)

OpenStudy (kainui):

Check this out: w'=v, v'=u, u'=w so let's differentiate: \[\large u^3+v^3+w^3 = 3uvw+1 \\ \large 3u^2 u' + 3v^2 v' + 3w^2 w' = 3u'vw + 3 uv'w + 3 uvw' \\ \large 3(u^2w+v^2 u+w^2v) = 3(wuw + uuw+ uvv)\]

OpenStudy (kainui):

I might have typed something wrong I don't know, there was a lot of stuff going on haha.

OpenStudy (freckles):

yes everything matches up except the w^2v=w^2u thing

ganeshie8 (ganeshie8):

*

OpenStudy (freckles):

I was trying it too...to see..I keep making a mistake I think...

OpenStudy (kainui):

I just worked it out on my paper, and it worked, the mistake is in the first term on the second half f the last equation, it should be w^2v, it's just the u and v look identical in that font lol. \[\large u^3+v^3+w^3 = 3uvw+1 \\ \large 3u^2 u' + 3v^2 v' + 3w^2 w' = 3u'vw + 3 uv'w + 3 uvw' \\ \large 3(u^2w+v^2 u+w^2v) = 3(wuw + uuw+ uvv)\]

OpenStudy (freckles):

my v's and u's look the same in font and in my handwriting

OpenStudy (freckles):

omg @Kainui that is so freaking awesome I thought this problem was going to have such terrible long solution

OpenStudy (kainui):

\[\Large v u \nu \mu\] Like all the same letter almost rofl. Haha yeah, I hate these types of problems honestly, so I always try to find an easy way out. =P

OpenStudy (freckles):

the question about this question is the power series mentioned by @ganeshie8 last question weird it talks about the Fibonacci numbers being coefficients ...blah blah black sheep. If I was smart I could have cheated if I only had looked at my book right next to me.

OpenStudy (kainui):

Although I don't know if the question is completely done, when we took the derivative we lost that +1 term, I don't know if that matters or what exactly.

OpenStudy (freckles):

the question before this question*

OpenStudy (kainui):

No, this question lol.

OpenStudy (freckles):

sometimes I don't realize what I'm typing until its too late

OpenStudy (kainui):

Like, what I'm trying to say is we didn't show it is true because for all we know: sin(x)=sin(x)+195 because they have the same derivative haha.

OpenStudy (freckles):

i have to type around a kitty all day and it is tough

OpenStudy (kainui):

Yeah my old roommate had a kitty several years ago, and it loved keyboards idk why haha.

OpenStudy (freckles):

\[(f(x))'=0\] f(x)=c so we need to show c is 1

OpenStudy (freckles):

u(0)=1 v(0)=0 w(0)=0

OpenStudy (kainui):

Awesome! Plug that in!

OpenStudy (freckles):

1+0+0-3(0)=c

OpenStudy (freckles):

1=c

OpenStudy (kainui):

Ok, now that was satisfyingly easier than I thought it'd be. Wow, done. That was fun.

OpenStudy (freckles):

I like the last bit because I thought to do it but now I will think to do the first part you mentioned hopefully later if I ever see a similar problem

OpenStudy (kainui):

Yeah I hope I do too. I wonder how often it will work for you though, since I don't think I've ever seen them defined like this before, kind of weird problem honestly.

OpenStudy (freckles):

well it was a weird problem it was in the problem plus section

OpenStudy (freckles):

so all of those problems are weird and fun

OpenStudy (freckles):

I will probably most likely never do this again though unless @xapproachesinfinity brings it up and @ganeshie8 says oh this is freckle's favorite problem :p

OpenStudy (freckles):

anyways goodnight!

ganeshie8 (ganeshie8):

lol don't burn the book, you might have to refer the same stuff again as xapproaches is following you like shadow :P gnite!

OpenStudy (kainui):

Haha well I think it's fun, these u,v, w functions are really the first power series I ever played with on my own outside of calculus 2 class. The thing that interested me was how e^x is its own first derivative, sinx is its own 4th derivative, and these are their own 3rd derivatives, etc... So I'll give you a challenge @freckles what is the power series representation of a function that is its own nth derivative?

OpenStudy (freckles):

and i guess you are looking for an answer other than e^x

OpenStudy (freckles):

I mean the power series representation for that function is: \[1+x+\frac{x^2}{2!}+\frac{x^3}{3!}+\frac{x^4}{4!}+\cdots \\ \text{ or in sigma notation } \\ \sum_{n=0}^{\infty} \frac{x^n}{n!}\]

OpenStudy (kainui):

Well, that is e^x, so maybe I should give some more examples of what I mean, e^-x and e^ix are their own 2nd and 4th derivatives respectively. \[\Large \frac{d}{dx} e^{-x} = -e^{-x} \\ \Large \frac{d^2}{dx^2} e^{-x} = e^{-x}\] \[\Large \frac{d}{dx} e^{ix} =ie^{ix} \\ \Large \frac{d^2}{dx^2} e^{ix} = -e^{ix} \\ \Large \frac{d^3}{dx^3} e^{ix} =-ie^{ix} \\ \Large \frac{d^4}{dx^4} e^{ix} = e^{ix}\] So what function satisfies this: \[\Large \frac{d^n}{dx^n} f(x)=f(x)\] Even though e^x satisfies this for all n, I'm looking for the function that satisfies this as little as possible... lol

OpenStudy (freckles):

\[(r-1)(r^{n-1}+r^{n-2}+r^{n-3}+\cdots+r^3+r^2+r+1) \\ \text{ so I need to solve: } r^{n-1}+r^{n-2}+\cdots +1=0\] ? So basically I need to solve this differential equation: \[y^{(n-1)}+y^{(n-2)}+ \cdots +y^{(1)}+y=0\] And I guess I could set my solution up as a power series... which I actually have already done and have gotten something not so great looking... But maybe I'm also choosing the wrong power series form...

OpenStudy (kainui):

Yeah, this is one way I think, but I think there's an easier way. Try to look at the examples I gave and see if you can find the common link between the examples I've given. I'll reveal the answer within an hour or two unless you really don't want me to tell you lol. =P

OpenStudy (freckles):

give me a day

OpenStudy (freckles):

post the answer in 24 hours or when you come back after 24 hours